# Towards Robust Model-Based Reinforcement Learning Against Adversarial Corruption

### v2

@@ 5.2 Analysis
Given an offline dataset $\mathcal{D}$, for any initial state $x^{1}=x\in\mathcal{X}$, the information coefficient with respect to space $-^\widehat{\mathcal{M}}$ is:
${\mathrm{IC}}^{\sigma}(\lambda,\widehat{\mathcal{M}},\mathcal{D})=\max_{h\in[H]}{\mathbb{E}}_{\pi_{*}}\bigg{[}\sup_{M,M^{\prime}\in\widehat{\mathcal{M}}}\frac{T\cdot l^{h}(M,M^{\prime},z^{h})^{2}/\sigma^{h}(z^{h})}{\lambda+\sum_{t=1}^{T}l_{t}^{h}(M,M^{\prime})^{2}/\sigma_{t}^{h}}\bigg{]},$
@@
To obtain an instance-independent upper bound, we follow Ye et al., 2023b to introduce a TV-norm-based version of the coverage condition for the dataset.
@@
$\frac{1}{T}\sum_{t=1}^{T}{\mathrm{TV}}(P_{M}^{h}(\cdot|z_{t}^{h})\|P_{M^{\prime}}^{h}(\cdot|z_{t}^{h}))^{2}\geq\mathrm{Cov}(\mathcal{M},\mathcal{D})\big{(}\rho^{h}(M,M^{\prime})\big{)}^{2}.$ (6)
This is a generally adopted condition in offline literature (Duan et al.,, 2020; Wang et al.,, 2020; Xiong et al.,, 2022; Di et al.,, 2023). Intuitively, this condition depicts that the dataset explores each direction of the space. In linear case where $\mathcal{M}^{h}=\{\langle\nu^{h}(M,x^{h+1}),\phi^{h}(\cdot)\rangle:\mathcal{X}\rightarrow{\mathbb{R}}\}$, according to Lemma D.1, this assumption is implied by the condition that the covariance matrix $T^{-1}\sum_{t=1}^{T}\phi(z_{t}^{h})\phi(z_{t}^{h})^{\top}$ is strictly positive definite, and $\mathrm{Cov}(\mathcal{M},\mathcal{D})$ has a $\Theta(d^{-1})$ dependence on $d$. Under this condition, the suboptimality upper bound yielded by Algorithm 2 is guaranteed.
@@
Supposing that Assumption 3.1 and 5.1 hold, if we choose $\lambda=\log|\mathcal{M}|$, $\alpha=\sqrt{\log|\mathcal{M}|\log^{2}B}/C$ and
@@
For a given corruption level $C$, data coverage coefficient $\mathrm{Cov}(\mathcal{M},\mathcal{D})$, dimension $d>3$ and episode length $H>2$, it the data size $T$ satisfied $T>\Omega\big{(}\mathrm{Cov}(\mathcal{M},\mathcal{D})/C\big{)}$, there exist an hard to learn instance such that the suboptimality for any algorithm is lower bounded by
@@
Here, our offline lower bound in Theorem 5.3 matches the the corruption term $CH/\big{(}T\mathrm{Cov}(\mathcal{M},\mathcal{D})\big{)}$ in our upper bound, up to a factor of $H$. This result demonstrates that our algorithm achieves near optimal suboptimality guarantee for defending against potential adversarial attacks.
@@ 6 Conclusion and Future Work
We delve into adversarial corruption for model-based RL, encompassing both online and offline settings. Our approach involves quantifying corruption by summing the differences between true and corrupted transitions. In the online setting, our algorithm, CR-OMLE, combines optimism and weighted log-likelihood principles, with weights employing a TV-norm-based uncertainty. Our analysis yields a $O(\sqrt{T}+C)$ regret upper bound for CR-OMLE and establishes a new lower bound for MDPS with transition corruption, aligning in terms of corruption-dependent terms. In the offline realm, we present CR-PMLE, a fusion of pessimism and MLE with uncertainty weights, offering theoretical assurances through instance-dependent and instance-independent upper bounds. Notably, the instance-independent bound necessitates a uniform coverage condition, where the corruption-dependent term nearly matches the lower bound. We anticipate that our findings will contribute theoretical insights to address practical challenges in model-based RL, particularly in the face of adversarial corruption.
Several future work are worth exploring: (1) Extending the uncertainty weighting technique to representation learning problems in RL (Jiang et al.,, 2017; Liu et al.,, 2022) would be an intriguing direction; and (2) In situations where the corruption level $C$ is unknown, our method relies on an optimistic estimation of corruption denoted by $\bar{C}$. There is a need for further investigation into the applicability of the model selection technique introduced by Wei et al., (2022) to the realm of model-based RL.
7 Impact Statements
This paper introduces research aimed at enhancing the resilience of model-based reinforcement learning (RL) when confronted with adversarial corruption and malicious attacks. The enhanced robustness achieved by this work not only enhances the reliability of RL systems but also contributes to the development of more secure and trustworthy AI technologies. It will also promote the greater public trust in AI applications, thereby promoting their responsible and ethical deployment across various domains.
@@ Appendix A Proof of Online Setting
We demonstrate the relationship between the weighted IR defined in (5) and eluder dimension in Definition 3.2. For simplicity, we consider a single step in the following lemma.
@@ Appendix B Proof of Offline Setting
First of all, we demonstrate that for the uncertainty weight iteration in Algorithm 3 converges and the output solution approximates the real uncertainty quantity. Since the convergence has been illustrated in Ye et al., 2023b , we restate the result in the following lemma.
Lemma B.1 (Lemma 3.1 of Ye et al., 2023a ).
@@
we have with probability at least $1-\delta$, for all $h\in[H]$ and all $t\in[T]$, $-^M_{*}\in\widehat{\mathcal{M}}$ and
@@ B.3 Offline Lower bound
$\rho^{h}(M,M^{\prime})^{2}$ $\leq\sum_{a\in\{a_{1},...,a_{d-1}\}}{\mathrm{TV}}(P^{h}_{M}(\cdot|x_{1},a)\|P^{h}_{M^{\prime}}(\cdot|x_{1},a))$
@@ C.1 Lemmas for Online Setting
To condense notations, we use the notation $l(M,\bar{M}_{t},z)={\mathrm{TV}}(P_{M}(\cdot|z)\|P_{\bar{M}_{t}}(\cdot|z))$. Now, we follow the three steps in the proof of Lemma 5.1 from Ye et al., 2023a .
@@ C.2 Lemmas for Offline Setting
The proof adapts the analysis of Lemma 4.1 in Ye et al., 2023b .
@@ D.1 Bounding TV-Eluder Dimension for Tabular MDPs
According to Lemma G.5 of Wang et al., (2023), we obtain that
@@ D.2 Bounding TV-Eluder Dimension for Linear MDPs
Then, by applying Lemma G.5 of Wang et al., (2023), we complete the proof. ∎
@@
Now, we use the linear MDP to illustrate the condition in Assumption 5.1. In the following lemma, we demonstrate that the condition holds as long as the learner has excess to a well-explored dataset (29), which is a wildly-adopted assumption in the literature of offline linear MDPs (Duan et al.,, 2020; Wang et al.,, 2020; Zhong et al., 2022a, ).
@@ Appendix E Technical Lemmas
Lemma E.1 (Azuma–Hoeffding inequality, Cesa-Bianchi and Lugosi, 2006).
@@
Lemma E.2 (Theorem 13.2 of Zhang, (2023)).
Consider a sequence of random functions $C_{1}(\mathcal{S}_{1}),\ldots,C_{t}(\mathcal{S}_{t}),\ldots$, with filtration $\{\mathcal{F}_{t}\}$ and $\mathcal{S}_{t}=(Z_{1},\ldots,Z_{t})$. We have for any $\delta\in(0,1)$ and $\lambda>0$:
${\mathbb{P}}\bigg{[}\exists n>0:-\sum_{i=1}^{n}C_{i}\leq\frac{\log(1/\delta)}{\lambda}+\frac{1}{\lambda}\sum_{i=1}^{n}\log{\mathbb{E}}_{Z_{i}^{(y)}}\exp(-\lambda C_{i})\bigg{]}\leq\delta,$
where $Z_{t}=(Z_{t}^{(x)},Z_{t}^{(y)})$. This implies that for all $\sigma>0$, with probability at least $1-\delta$:
[…]
Lemma E.3 (Theorem 13.6 of Zhang, (2023)).
Consider a sequence of random functions $C_{1}(\mathcal{S}_{1}),\ldots,C_{t}(\mathcal{S}_{t})$. Assume that $\xi_{i}\geq{\mathbb{E}}_{Z_{i}^{(y)}}\xi_{i}-b$ for some constant $b>0$. Then for any $\lambda\in(0,3/b)$, with probability at least $1-\delta$:
@@
Lemma E.5 (Proposition B.11 of Zhang, (2023)).
@@
Lemma E.6 (Elliptical Potential Lemma (Dani et al.,, 2008; Rusmevichientong and Tsitsiklis,, 2010; Abbasi-Yadkori et al.,, 2011)).

### v3

@@ 5.2 Analysis
Given an offline dataset $\mathcal{D}$, for any initial state $x^{1}=x\in\mathcal{X}$, the information coefficient with respect to space $\widehat{\mathcal{M}}$ is:
$\mathrm{IC}^{\sigma}(\lambda,\widehat{\mathcal{M}},\mathcal{D})$
$=\max_{h\in[H]}{\mathbb{E}}_{\pi_{*}}\bigg{[}\sup_{M,M^{\prime}\in\widehat{\mathcal{M}}}\frac{T\cdot l^{h}(M,M^{\prime},z^{h})^{2}/\sigma^{h}(z^{h})}{\lambda+\sum_{t=1}^{T}l_{t}^{h}(M,M^{\prime})^{2}/\sigma_{t}^{h}}\bigg{]},$
@@
To obtain an instance-independent upper bound, we follow Ye et al. (2023b) to introduce a TV-norm-based version of the coverage condition for the dataset.
@@
$\frac{1}{T}\sum_{t=1}^{T}{\mathrm{TV}}(P_{M}^{h}(\cdot|z_{t}^{h})\|P_{M^{\prime}}^{h}(\cdot|z_{t}^{h}))^{2}\geq\mathrm{Cov}(\mathcal{M},\mathcal{D})\big{(}\rho^{h}(M,M^{\prime})\big{)}^{2}.$
This is a generally adopted condition in offline literature (Duan et al., 2020; Wang et al., 2020; Xiong et al., 2022; Di et al., 2023). Intuitively, this condition depicts that the dataset explores each direction of the space. In linear case where $\mathcal{M}^{h}=\{\langle\nu^{h}(M,x^{h+1}),\phi^{h}(\cdot)\rangle:\mathcal{X}\rightarrow{\mathbb{R}}\}$, according to Lemma D.4, this assumption is implied by the condition that the covariance matrix $T^{-1}\sum_{t=1}^{T}\phi(z_{t}^{h})\phi(z_{t}^{h})^{\top}$ is strictly positive definite, and $\mathrm{Cov}(\mathcal{M},\mathcal{D})$ has a $\Theta(d^{-1})$ dependence on $d$. Under this condition, the suboptimality upper bound yielded by Algorithm 2 is guaranteed.
@@
Supposing that Assumption 3.1 and 5.3 hold, if we choose $\lambda=\log|\mathcal{M}|$, $\alpha=\sqrt{\log|\mathcal{M}|\log^{2}B}/C$ and
@@
Remark 5.5 (Comparison between Online and Offline Results).
Both of the online and offline bounds are composed of the main term (uncorrupted error) and the corruption error, and the corruption error is sublinear as long as the corruption level $C$ is sublinear w.r.t sample size $T$. The main difference is that the online bound is affected by the eluder dimension and the offline bound is affected by the covering coefficient. This difference comes from the ability to continue interacting with the environment. In the online setting, the agent can explore the model space extensively, thus, the final regret bound is controlled by the complexity of the model space. In contrast, exploration in the offline setting is constrained by a given dataset, thus the final bound is determined by the coverage quality of the dataset.
@@
For a given corruption level $C$, data coverage coefficient $\mathrm{Cov}(\mathcal{M},\mathcal{D})$, dimension $d>3$ and episode length $H>2$, it the data size $T$ satisfied $T>\Omega\big{(}\mathrm{Cov}(\mathcal{M},\mathcal{D})/C\big{)}$, there exist a hard to learn instance such that the suboptimality for any algorithm is lower bounded by
@@
The offline lower bound in Theorem 5.6 matches the corruption term $CH/\big{(}T\mathrm{Cov}(\mathcal{M},\mathcal{D})\big{)}$ in our upper bound, up to a factor of $H$. This result demonstrates that our algorithm achieves near-optimal suboptimality guarantee for defending against adversarial attacks.
@@ 6 Conclusion and Future Work
We delve into adversarial corruption for model-based RL, encompassing both online and offline settings. Our approach involves quantifying corruption by summing the differences between true and corrupted transitions. In the online setting, our algorithm, CR-OMLE, combines optimism and weighted log-likelihood principles, with weights employing a TV-norm-based uncertainty. Our analysis yields a $O(\sqrt{T}+C)$ regret upper bound for CR-OMLE and establishes a new lower bound for MDPS with transition corruption, aligning in terms of corruption-dependent terms. In the offline realm, we present CR-PMLE, a fusion of pessimism and MLE with uncertainty weights, offering theoretical assurances through instance-dependent and instance-independent upper bounds. Notably, the instance-independent bound necessitates a uniform coverage condition, where the corruption-dependent term nearly matches the lower bound. We anticipate that our findings will contribute theoretical insights to address practical challenges in model-based RL, particularly for adversarial corruption.
Several future works are worth exploring: (1) Extending the uncertainty weighting technique to representation learning problems in RL (Jiang et al., 2017; Liu et al., 2022) would be an intriguing direction; and (2) In situations where the corruption level $C$ is unknown, our method relies on an optimistic estimation of corruption denoted by $\bar{C}$. There is a need for further investigation into the applicability of the model selection technique introduced by Wei et al. (2022) to the realm of model-based RL.
Acknowledgements
The authors would like to thank the anonymous reviewers for many insightful comments and suggestions. JH and QG are supported by the National Science Foundation CAREER Award 1906169 and research fund from UCLA-Amazon Science Hub. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
Impact Statement
This paper introduces research aimed at enhancing the resilience of model-based reinforcement learning (RL) when confronted with adversarial corruption and malicious attacks. The enhanced robustness achieved by this work not only enhances the reliability of RL systems but also contributes to the development of more secure and trustworthy AI technologies. It will also promote greater public trust in AI applications, thereby promoting their responsible and ethical deployment across various domains.
@@ Appendix A Proof of Online Setting
We demonstrate the relationship between the weighted IR defined in (4.4) and eluder dimension in Definition 3.3. For simplicity, we consider a single step in the following lemma.
@@ Appendix B Proof of Offline Setting
First of all, we demonstrate that for the uncertainty weight iteration in Algorithm 3 converges and the output solution approximates the real uncertainty quantity. Since the convergence has been illustrated in Ye et al. (2023b), we restate the result in the following lemma.
Lemma B.1 (Lemma 3.1 of Ye et al. (2023a)).
@@
we have with probability at least $1-\delta$, for all $h\in[H]$ and all $t\in[T]$, $M_{*}\in\widehat{\mathcal{M}}$ and
@@ B.3 Offline Lower bound
$\rho^{h}(M,M^{\prime})^{2}$ $\leq\sum_{a\in\{a_{1},...,a_{d-1}\}}{\mathrm{TV}}(P^{h}_{M}(\cdot|x_{1},a)\|P^{h}_{M^{\prime}}(\cdot|x_{1},a))^{2}$
@@ C.1 Lemmas for Online Setting
To condense notations, we use the notation $l(M,\bar{M}_{t},z)={\mathrm{TV}}(P_{M}(\cdot|z)\|P_{\bar{M}_{t}}(\cdot|z))$. Now, we follow the three steps in the proof of Lemma 5.1 from Ye et al. (2023a).
@@ C.2 Lemmas for Offline Setting
The proof adapts the analysis of Lemma 4.1 in Ye et al. (2023b).
@@ D.1 Bounding TV-Eluder Dimension for Tabular MDPs
According to Lemma G.5 of Wang et al. (2023), we obtain that
@@ D.2 Bounding TV-Eluder Dimension for Linear MDPs
Then, by applying Lemma G.5 of Wang et al. (2023), we complete the proof. ∎
@@
Now, we use the linear MDP to illustrate the condition in Assumption 5.3. In the following lemma, we demonstrate that the condition holds as long as the learner has excess to a well-explored dataset (28), which is a wildly-adopted assumption in the literature of offline linear MDPs (Duan et al., 2020; Wang et al., 2020; Zhong et al., 2022a).
@@ Appendix E Technical Lemmas
Lemma E.1 (Azuma–Hoeffding inequality, Cesa-Bianchi & Lugosi 2006).
@@
Lemma E.2 (Theorem 13.6 of Zhang (2023)).
Consider a sequence of random functions $C_{1}(\mathcal{S}_{1}),\ldots,C_{t}(\mathcal{S}_{t})$. Assume that $\xi_{i}\leq{\mathbb{E}}_{Z_{i}^{(y)}}\xi_{i}+b$ for some constant $b>0$. Then for any $\lambda\in(0,3/b)$, with probability at least $1-\delta$:
@@
Lemma E.4 (Proposition B.11 of Zhang (2023)).
@@
Lemma E.5 (Elliptical Potential Lemma (Dani et al., 2008; Rusmevichientong & Tsitsiklis, 2010; Abbasi-Yadkori et al., 2011)).